OpenStudy (anonymous):

Find the point on the line 6x+y=9 that is closest to the point (8−9).

OpenStudy (anonymous):

the closest point on the line 6x+y=9 is the point of intersection of the line and the line perpendicular to it passing through the point (8,-9). any line perpendicular to 6x+y=9 is of the form x-6y=c in this particular case the perpendicular line passes through the point (8,-9) so the equation of the perpendicular line is x-6y=62 now all you have to do is find out the point of intersection between 6x+y=9 and x-6y=62 from equation i, 36x+6y=54 from equation ii, x-6y=62 adding those two, 37x=116 or, x=116/37 so, y=9-6*116/37

OpenStudy (anonymous):

Thanks :), I got it, although you make it look easier lol